# ESGSenticNet: A Neurosymbolic Knowledge Base for Corporate Sustainability Analysis

### v1

@@ Abstract
Evaluating corporate sustainability performance is essential to drive sustainable business practices, amid the need for a more sustainable economy. However, this is hindered by the complexity and volume of corporate sustainability data (i.e. sustainability disclosures), not least by the effectiveness of the NLP tools used to analyse them. To this end, we identify three primary challenges – immateriality, complexity, and subjectivity, that exacerbate the difficulty of extracting insights from sustainability disclosures. To address these issues, we introduce ESGSenticNet, a publicly available knowledge base for sustainability analysis. ESGSenticNet is constructed from a neurosymbolic framework that integrates specialised concept parsing, GPT-4o inference, and semi-supervised label propagation, together with a hierarchical taxonomy. This approach culminates in a structured knowledge base of 44k knowledge triplets – (‘halve carbon emission’, supports, ‘emissions control’), for effective sustainability analysis. Experiments indicate that ESGSenticNet, when deployed as a lexical method, more effectively captures relevant and actionable sustainability information from sustainability disclosures compared to state of the art baselines. Besides capturing a high number of unique ESG topic terms, ESGSenticNet outperforms baselines on the ESG relatedness and ESG action orientation of these terms by 26% and 31% respectively. These metrics describe the extent to which topic terms are related to ESG, and depict an action toward ESG. Moreover, when deployed as a lexical method, ESGSenticNet does not require any training, possessing a key advantage in its simplicity for non-technical stakeholders.
@@
journal: Information Processing and Management
@@
[8]organization=Honorary Research Fellow, Imperial College London,

### v2

@@ Abstract
Identifying Environmental, Social, and Governance (ESG) topics in sustainability reports, or ESG topic analysis, is essential for assessing a company’s ESG performance. While commonly used for this task, existing NLP approaches face three key challenges: irrelevance, the extracted topics are unrelated to ESG, immateriality, the extracted topics lack meaningful value, and limited organisation, the extracted topics have poor alignment with ESG frameworks.
To address these challenges, we introduce ESGSenticNet, a publicly available knowledge base built from 1,679 SGX sustainability reports. ESGSenticNet is constructed using a neurosymbolic framework—combining symbolic reasoning (structured parsing rules) with neural language models (GPT-4o)—to identify and categorise ESG-related concepts. The process begins by extracting concepts (or meaningful phrases) from reports using parsing methods, which are then labelled according to a hierarchical ESG taxonomy using GPT-4o. These labelled concepts are further expanded using a graph-based method that propagates labels to similar concepts without requiring additional manual annotation. The final knowledge base consists of 44,232 structured entries, or “triples,” in the format (concept, relation, ESG topic)—e.g., (“halve carbon emission”, supports, “emissions control”)—which serve as ESG lexicons for topic analysis. This allows ESG topics to be identified by matching concepts found in reports to their associated topics.
We evaluate ESGSenticNet on 319 sustainability reports against existing NLP methods. ESGSenticNet outperforms baselines, producing topic terms with 76% ESG-relatedness (+26% improvement) and 84% ESG action-orientation (+34% improvement), while capturing a high number of unique ESG terms (359). Human evaluation of 500 concepts shows that ESGSenticNet achieves 86%+ accuracy in categorising concepts under correct ESG topics. This suggests that during topic analysis, ESGSenticNet shows a strong degree of reliability in aligning topic terms with their intended topics. Additionally, ESGSenticNet’s deployment requires no training, making it accessible for stakeholders regardless of resources or expertise. Our key contributions include: (1) developing a neurosymbolic framework for deriving sustainability lexicons, which can be potentially extended to other contexts such as climate risk disclosures; (2) introducing new metrics for ESG topic analysis, which we argue are more meaningful for the task at hand; and (3) releasing ESGSenticNet as a public tool that effectively addresses the core challenges of irrelevance, immateriality, and limited organisation in ESG topic analysis.
@@
journal: npj Artificial Intelligence
@@
[8]organization=Honorary Research Fellow, Imperial College London, addressline=Exhibition Road, London SW7 2AZ, country=United Kingdom